# references.bib

### v3

@@ 4.2. Proof of the cohomological integrality theorem
The injectivity is Proposition 4.26. The surjectivity comes from Lemmas 4.16 and 2.15 and the formula
$\mathrm{Ind}_{\mu}(f)=\frac{1}{\#W^{\mu}}\sum_{\overline{w}\in W/W_{\mu}}w^{\prime}\cdot\left(\sum_{w\in W_{\mu}}\varepsilon_{V,\mu}(w)w\cdot f\right)(w^{\prime}\cdot k_{\mu})$
deduced from Propositions 2.9 and 4.8 which implies, using Proposition 4.10, that $\mathrm{Ind}_{\mu}(f)$ vanishes if $f$ is not in the $\varepsilon_{V,\mu}$-isotypic component of $\mathcal{P}_{\mu}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\mu})$ for the $W_{\mu}$-action. ∎
@@ 4.3. The BPS subspaces
In analogy to the case of quivers [efimov2012cohomological, meinhardt2019donaldson], we may call the subspaces $\mathcal{P}_{\lambda}\subseteq\mathrm{H}^{*}(V^{\lambda}/G^{\lambda})$ the BPS subspaces. Although we lack both qualitative and quantitative properties of these spaces, there are cases when we can prove it vanishes.
@@
Let $\lambda\in X_{*}(T)$. If $\ker(G^{\lambda}\rightarrow\mathrm{GL}(V^{\lambda}))^{0}\not\subseteq Z(G^{\lambda})$, then $\mathcal{P}_{\lambda}=\{0\}$.
@@
By possibly replacing $V$ by $V^{\lambda}$ and $G$ by $G^{\lambda}$, we may assume that $\lambda$ is the trivial cocharacter of $T$. If $\ker(G\rightarrow\mathrm{GL}(V))\not\subseteq Z(G)$, then there exists a cocharacter $\mu\colon\mathbf{G}_{\mathrm{m}}\rightarrow\ker(G\rightarrow\mathrm{GL}(V))$ which does not factor through $Z(G)$. By conjugating $\mu$ by some element $g\in G$, we may assume that $\mu\colon\mathbf{G}_{\mathrm{m}}\rightarrow T\cap\ker(G\rightarrow\mathrm{GL}(V))$. We have $\overline{\mu}\in\mathscr{P}_{V}\smallsetminus\{\overline{0}\}$. Then, the induction diagram reads
@@
and the induction map $\mathrm{Ind}_{\mu}\colon\mathrm{H}^{*}(V/G^{\mu})\rightarrow\mathrm{H}^{*}(V/G)$ is surjective. Hence, $\mathcal{P}_{0}=\{0\}$. ∎
5. Vanishing bounds
In this section, we prove Proposition 1.4.
First, we have an inclusion $\mathcal{P}_{\lambda}^{k}\subseteq\mathcal{H}_{\lambda}^{k}=\mathrm{H}^{k+d_{\lambda}}(V^{\lambda}/G^{\lambda})$, and the latter vector space is zero for $k<-d_{\lambda}=\dim G^{\lambda}-\dim V^{\lambda}$. The proof of the upper bound uses [efimov2012cohomological, Lemma 3.6] and is analogous to the proof of [efimov2012cohomological, Theorem 3.5]. Namely, [efimov2012cohomological, Lemma 3.6] gives $(\mathcal{A}_{\lambda}^{\mathrm{prim}})^{d}\subseteq J_{\lambda}$ for $d\geqslant\dim V^{\lambda}$ (for the unshifted cohomological gradings), where $(\mathcal{A}_{\lambda}^{\mathrm{prim}})^{d}$ denotes degree $d$ polynomials inside $\mathcal{A}_{\lambda}^{\mathrm{prim}}$.
@@ 6. Examples
In this section, we give some explicit examples of cohomological integrality isomorphisms for some choices of pairs $(G,V)$ of a reductive group $G$ and a symmetric representation $V$ of $G$. We give each time the explicit formula for the cohomological integrality isomorphisms. In each case, it is possible to verify by hand that they are indeed isomorphisms.
@@ 6.1. $(\mathbf{C}^{*})^{2}\curvearrowright\mathbf{C}^{2}\oplus(\mathbf{C}^{2})^{\vee}$
We have $\mathscr{P}_{V}=\{\overline{\lambda_{0}},\overline{\lambda_{1}},\overline{\lambda_{2}},\overline{\lambda_{3}}\}$ where
$\begin{matrix}\lambda_{1}&\colon&\mathbf{G}_{\mathrm{m}}&\rightarrow&(\mathbf{C}^{*})^{2}\\ &&v&\mapsto&(v,1)\end{matrix}$
$\begin{matrix}\lambda_{2}&\colon&\mathbf{G}_{\mathrm{m}}&\rightarrow&(\mathbf{C}^{*})^{2}\\ &&v&\mapsto&(1,v)\end{matrix}$
$\begin{matrix}\lambda_{3}&\colon&\mathbf{G}_{\mathrm{m}}&\rightarrow&(\mathbf{C}^{*})^{2}\\ &&v&\mapsto&(v,v)\end{matrix}$
We have $V^{\lambda_{0}}=V$, $V^{\lambda_{1}}=(0\oplus\mathbf{C})\oplus(0\oplus\mathbf{C}^{\vee})$, $V^{\lambda_{2}}=(\mathbf{C}\oplus 0)\oplus(\mathbf{C}^{\vee}\oplus 0)$ and $V^{\lambda_{3}}=\{0\}$. We can describe the induction maps as follows:
$\begin{matrix}\mathrm{Ind}_{\lambda_{1},\lambda_{0}}&\colon&\mathbf{Q}[x_{1},x_{2}]&\rightarrow&\mathbf{Q}[x_{1},x_{2}]\\ &&f&\mapsto&x_{1}f\end{matrix}$
$\begin{matrix}\mathrm{Ind}_{\lambda_{2},\lambda_{0}}&\colon&\mathbf{Q}[x_{1},x_{2}]&\rightarrow&\mathbf{Q}[x_{1},x_{2}]\\ &&f&\mapsto&x_{2}f\end{matrix}$
$\begin{matrix}\mathrm{Ind}_{\lambda_{3},\lambda_{0}}&\colon&\mathbf{Q}[x_{1},x_{2}]&\rightarrow&\mathbf{Q}[x_{1},x_{2}]\\ &&f&\mapsto&x_{1}x_{2}f\end{matrix}$
$\begin{matrix}\mathrm{Ind}_{\lambda_{3},\lambda_{1}}&\colon&\mathbf{Q}[x_{1},x_{2}]&\rightarrow&\mathbf{Q}[x_{1},x_{2}]\\ &&f&\mapsto&x_{2}f\end{matrix}$
$\begin{matrix}\mathrm{Ind}_{\lambda_{3},\lambda_{2}}&\colon&\mathbf{Q}[x_{1},x_{2}]&\rightarrow&\mathbf{Q}[x_{1},x_{2}]\\ &&f&\mapsto&x_{1}f\end{matrix}$
@@
$\mathcal{P}_{\lambda_{0}}=\mathbf{Q}\subseteq\mathbf{Q}[x_{1},x_{2}],\quad\mathcal{P}_{\lambda_{1}}=\mathbf{Q},\quad\mathcal{P}_{\lambda_{2}}=\mathbf{Q},\quad\mathcal{P}_{\lambda_{3}}=\mathbf{Q}$
and the cohomological integrality isomorphism is
$\begin{matrix}\mathbf{Q}\oplus\mathbf{Q}[x_{1}]\oplus\mathbf{Q}[x_{2}]\oplus\mathbf{Q}[x_{1},x_{2}]&\rightarrow&\mathbf{Q}[x_{1},x_{2}]\\ (a_{0},f(x_{1}),g(x_{2}),h(x_{1},x_{2}))&\mapsto&a_{0}+x_{1}f(x_{1})+x_{2}g(x_{2})+x_{1}x_{2}h(x_{1},x_{2}).\end{matrix}$
@@ 6.2. $\mathrm{GL}_{2}(\mathbf{C})\curvearrowright\mathbf{C}^{2}\oplus(\mathbf{C}^{2})^{\vee}$
where the superscript $t$ indicates the transpose. We let $T\coloneqq(\mathbf{C}^{*})^{2}$ be the standard torus. The Weyl group is $W=\mathfrak{S}_{2}$. We have $\mathscr{P}_{V}/W=\{\widetilde{\lambda_{0}},\widetilde{\lambda_{1}},\widetilde{\lambda_{2}}\}$ where
@@
$\begin{matrix}\lambda_{1}&\colon&\mathbf{G}_{\mathrm{m}}&\rightarrow&(\mathbf{C}^{*})^{2}\\ &&v&\mapsto&(v,1)\end{matrix}$
$\begin{matrix}\lambda_{2}&\colon&\mathbf{G}_{\mathrm{m}}&\rightarrow&(\mathbf{C}^{*})^{2}\\ &&v&\mapsto&(v,v)\end{matrix}$
Therefore, we have $V^{\lambda_{0}}=V$, $V^{\lambda_{1}}=(0\oplus\mathbf{C})\oplus(0\oplus\mathbf{C}^{\vee})$ and $V^{\lambda_{2}}=\{0\}$. We can describe the induction maps as follows:
$\begin{matrix}\mathrm{Ind}_{\lambda_{1},\lambda_{0}}&\colon&\mathbf{Q}[x_{1},x_{2}]&\rightarrow&\mathbf{Q}[x_{1},x_{2}]^{\mathfrak{S}_{2}}=\mathbf{Q}[x_{1}+x_{2},x_{1}x_{2}]\\ &&f&\mapsto&\frac{x_{1}}{x_{1}-x_{2}}f(x_{1},x_{2})+\frac{x_{2}}{x_{2}-x_{1}}f(x_{2},x_{1})\end{matrix}$
$\begin{matrix}\mathrm{Ind}_{\lambda_{2},\lambda_{0}}&\colon&\mathbf{Q}[x_{1},x_{2}]&\rightarrow&\mathbf{Q}[x_{1},x_{2}]^{\mathfrak{S}_{2}}=\mathbf{Q}[x_{1}+x_{2},x_{1}x_{2}]\\ &&f&\mapsto&\frac{x_{1}x_{2}}{x_{1}-x_{2}}f(x_{1},x_{2})+\frac{x_{1}x_{2}}{x_{2}-x_{1}}f(x_{2},x_{1})\end{matrix}$
$\begin{matrix}\mathrm{Ind}_{\lambda_{2},\lambda_{1}}&\colon&\mathbf{Q}[x_{1},x_{2}]&\rightarrow&\mathbf{Q}[x_{1},x_{2}]\\ &&f&\mapsto&x_{2}f\end{matrix}$
@@
A direct sum complement of the image of $\mathrm{Ind}_{\lambda_{2},\lambda_{1}}$ is $\mathbf{Q}[x_{1}]=\mathcal{P}_{\lambda_{1}}\otimes\mathbf{Q}[x_{1}]$.
Therefore, we have $\mathcal{P}_{\lambda_{0}}=0$, $\mathcal{P}_{\lambda_{1}}=\mathbf{Q}$, $\mathcal{P}_{\lambda_{2}}=\mathbf{Q}$ and the cohomological integrality isomorphism reads
$\begin{matrix}\mathbf{Q}[x_{1}]\oplus\mathbf{Q}[x_{1},x_{2}]^{\mathrm{sgn}}&\rightarrow&\mathbf{Q}[x_{1},x_{2}]^{\mathfrak{S}_{2}}\\ (f(x_{1}),g(x_{1},x_{2}))&\mapsto&\frac{x_{1}f(x_{1})-x_{2}f(x_{2})}{x_{1}-x_{2}}+2x_{1}x_{2}\frac{g(x_{1},x_{2})}{x_{1}-x_{2}}\end{matrix}$
where $\mathbf{Q}[x_{1},x_{2}]^{\mathrm{sgn}}$ is the sign-isotypic component, since the kernel $\frac{x_{1}x_{2}}{x_{1}-x_{2}}$ changes sign when we exchange the variables $x_{1},x_{2}$.

### v6

@@ 4.2. Proof of the cohomological integrality theorem
The injectivity is Proposition 4.23. The surjectivity comes from Lemmas 4.14 and 2.17. Indeed, by Lemma 4.14, any element $f\in\mathrm{H}^{*}(V/G)$ has the form $f=\sum_{\mu\preceq 0}\mathrm{Ind}_{\mu}(g_{\mu})$ for some $g_{\mu}\in\mathcal{P}_{\mu}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\mu})$. By Lemma 2.17, we may rewrite $f=\sum_{\widetilde{\mu}\in\mathscr{P}_{V}/W}\mathrm{Ind}_{\mu}(\widetilde{g}_{\mu})$. We have the formula
$\mathrm{Ind}_{\mu}(\widetilde{g}_{\mu})=\frac{1}{\#W^{\mu}}\sum_{\overline{w^{\prime}}\in W/W_{\mu}}w^{\prime}\cdot\left(\sum_{w\in W_{\mu}}\varepsilon_{V,\mu}(w)^{-1}w\cdot\widetilde{g}_{\mu}\right)(w^{\prime}\cdot k_{\mu})$
for $\widetilde{g}_{\mu}\in\mathcal{H}_{\mu}$, deduced from Propositions 2.11 and 4.6 which implies, using Proposition 4.8, that $\mathrm{Ind}_{\mu}(\widetilde{g}_{\mu})$ vanishes if $f$ is not in the $\varepsilon_{V,\mu}$-isotypic component of $\mathcal{P}_{\mu}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\mu})$ for the $W_{\mu}$-action. Therefore, we may assume that $\widetilde{g}_{\mu}$ is $\varepsilon_{V,\lambda}$-invariant. This concludes the proof of the corollary. ∎
@@ 4.3. The BPS subspaces
In analogy to the case of quivers [Efi12, MR19] and also [DM20] for quivers with potential, we may call the subspaces $\mathcal{P}_{\lambda}\subseteq\mathrm{H}^{*}(V^{\lambda}/G^{\lambda})$ the BPS subspaces. Although we lack both qualitative and quantitative properties of these spaces, there are cases when we can prove they vanish.
@@
Let $\lambda\in X_{*}(T)$. If $\ker(G^{\lambda}\rightarrow\mathrm{GL}(V^{\lambda}))^{\circ}\not\subseteq Z(G^{\lambda})$, then $\mathcal{P}_{\lambda}=\{0\}$.
@@
By possibly replacing $V$ by $V^{\lambda}$ and $G$ by $G^{\lambda}$, we may assume that $\lambda$ is the trivial cocharacter of $T$. If $\ker(G\rightarrow\mathrm{GL}(V))^{\circ}\not\subseteq Z(G)$, then there exists a cocharacter $\mu\colon\mathbf{G}_{\mathrm{m}}\rightarrow\ker(G\rightarrow\mathrm{GL}(V))\subseteq T$ which does not factor through $Z(G)$. We have $\overline{\mu}\in\mathscr{P}_{V}\smallsetminus\{\overline{0}\}$. Then, the induction diagram reads
@@
and the induction map $\mathrm{Ind}_{\mu}\colon\mathrm{H}^{*}(V/G^{\mu})=\mathrm{H}^{*}(\mathrm{pt}/T)^{W^{\mu}}\rightarrow\mathrm{H}^{*}(V/G)=\mathrm{H}^{*}(\mathrm{pt}/T)$ is the map
$f\mapsto\sum_{w\in W/W^{\mu}}w\cdot\left(\frac{f}{\prod_{\alpha\in\mathcal{W}^{\mu<0}(\mathfrak{g})}\alpha}\right)\,.$
It is surjective. Hence, $\mathcal{P}_{0}=\{0\}$. ∎
5. Vanishing bounds for the BPS spaces
We recall that $\mathcal{H}_{\lambda}=\mathrm{H}^{*}(V^{\lambda}/G^{\lambda})$ for $\lambda\in X_{*}(T)$. We recall the chosen $\mathbf{Z}$-grading on $\mathcal{H}_{\lambda}$: $\mathcal{H}_{\lambda}^{k}=\mathrm{H}^{k+d_{\lambda}}(V^{\lambda}/G^{\lambda})$ for $k\in\mathbf{Z}$. This is just a shift of the cohomological grading, chosen so that induction morphisms preserve this grading (Lemma 2.7). In this section, we prove Proposition 1.5.
Proof of the lower bound in Proposition 1.5.
First, we have an inclusion $\mathcal{P}_{\lambda}^{k}\subseteq\mathcal{H}_{\lambda}^{k}=\mathrm{H}^{k+d_{\lambda}}(V^{\lambda}/G^{\lambda})$, and the latter vector space is zero for $k<-d_{\lambda}=\dim G^{\lambda}-\dim V^{\lambda}$. ∎
The proof of the upper bound uses [Efi12, Lemma 3.6] and is analogous to the proof of [Efi12, Theorem 3.5].
Lemma 5.1 ([Efi12, Lemma 3.6]).
Let $k$ be a field and let $B=k[z_{1},\ldots,z_{n}]$ be an algebra of polynomials with $\deg(z_{i})=2$ (to match the cohomological degrees). Suppose that $\ell_{1},\ldots,\ell_{s}\in B^{1}$ are pairwise linearly independent linear forms in $z_{i}$. Let $\{P_{1},\ldots,P_{r}\}\subseteq B$ be non-empty set of polynomials of the form
$P_{i}=\prod_{j=1}^{s}\ell_{j}^{d_{i,s}}$
for some $d_{i,j}\in\mathbf{Z}_{\geqslant 0}$. Let $d_{j}\coloneqq\max_{1\leqslant i\leqslant r}d_{i,j}$ for $1\leqslant j\leqslant s$. Then, the following are equivalent:
  1. (i)
    $B^{2d}\subseteq(P_{1},\ldots,P_{r})$ for $d\geqslant\sum_{j=1}^{s}d_{j}-n+1$,
  2. (ii)
    the ideal $(P_{1},\ldots,P_{r})\subseteq B$ has finite codimension,
  3. (iii)
    For any sequence $p_{1},\ldots,p_{r}$ of numbers in $\{1,\ldots,s\}$ such that $d_{i,p_{i}}>0$ for $1\leqslant i\leqslant r$, the linear forms generate the space $B^{1}$.
Proof of the upper bound in Proposition 1.5.
As we already noticed in the proof of Proposition 4.16, if we replace $V$ by $V\times\mathfrak{g}$, then we have inclusions $\mathcal{P}_{V,\lambda}\subseteq\mathcal{P}_{V\times\mathfrak{g},\lambda}$, and therefore, it suffices to give the upper bound for $\mathcal{P}_{V\times\mathfrak{g},\lambda}$. We set $B\coloneqq\mathrm{Sym}[(\mathfrak{t}/\mathfrak{g}_{\lambda})^{*}]=\mathcal{A}_{\lambda}^{\mathrm{prim}}$, the (polynomial) algebra of functions on $\mathfrak{g}/\mathfrak{g}_{\lambda}$. The proof of Proposition 4.16 shows that the ideal $J_{\lambda}\subseteq\mathrm{Sym}[(\mathfrak{t}/\mathfrak{g}_{\lambda})^{*}]$ has finite codimension, that is satisfies the second point of Lemma 5.1. In our case, the quantity $\sum_{j=1}^{s}d_{j}$ is equal to $\frac{\dim V^{\lambda}-\dim V^{T}}{2}$. Therefore, an upper bound for $\sum_{j=1}^{s}d_{j}-n+1$ is $\frac{\dim V^{\lambda}}{2}$.
Now, Lemma 5.1 gives $(\mathcal{A}_{\lambda}^{\mathrm{prim}})^{2d}\subseteq J_{\lambda}$ for $d\geqslant\frac{\dim V^{\lambda}}{2}$ (for the unshifted cohomological gradings), where $(\mathcal{A}_{\lambda}^{\mathrm{prim}})^{e}$ denotes degree $e$ polynomials inside $\mathcal{A}_{\lambda}^{\mathrm{prim}}$. By taking into account the cohomological shift by $\dim G^{\lambda}-\dim V^{\lambda}$, we obtain the upper bound: $\mathcal{P}_{\lambda}^{k}=0$ for $k>\dim V^{\lambda}+\dim G^{\lambda}-\dim V^{\lambda}=\dim G^{\lambda}$. ∎
@@ 6. Examples
In this section, we give some explicit examples of cohomological integrality isomorphisms for some choices of pairs $(G,V)$ of a reductive group $G$ and a (weakly) symmetric representation $V$ of $G$. We give each time the explicit formula for the cohomological integrality isomorphisms produced by Theorem 1.1. In each case, it is possible to verify by hand that they are indeed isomorphisms.
@@ 6.1. $(\mathbf{C}^{*})^{2}\curvearrowright\mathbf{C}^{2}\oplus(\mathbf{C}^{2})^{\vee}$
We can identify the set of equivalence classes of cocharacters (§2.6) $\mathscr{P}_{V}=\{\overline{\lambda_{0}},\overline{\lambda_{1}},\overline{\lambda_{2}},\overline{\lambda_{3}}\}$ where
[…]
$\begin{matrix}\lambda_{3}&\colon&\mathbf{G}_{\mathrm{m}}&\rightarrow&(\mathbf{C}^{*})^{2}\\ &&v&\mapsto&(v^{-1},v^{-1})\,.\end{matrix}$
We have $V^{\lambda_{0}}=V$, $V^{\lambda_{1}}=(0\oplus\mathbf{C})\oplus(0\oplus\mathbf{C}^{\vee})$, $V^{\lambda_{2}}=(\mathbf{C}\oplus 0)\oplus(\mathbf{C}^{\vee}\oplus 0)$ and $V^{\lambda_{3}}=\{0\}$. We have $G^{\lambda_{0}}=G^{\lambda_{1}}=G^{\lambda_{2}}=G^{\lambda_{3}}=(\mathbf{C}^{*})^{2}$. We can describe the induction maps as follows:
$\begin{matrix}\mathrm{Ind}_{\lambda_{1},\lambda_{0}}&\colon&\overbrace{\mathbf{Q}[x_{1},x_{2}]}^{=\mathrm{H}^{*}(V^{\lambda_{1}}/G^{\lambda_{1}})}&\rightarrow&\overbrace{\mathbf{Q}[x_{1},x_{2}]}^{=\mathrm{H}^{*}(V/G)}\\ &&f&\mapsto&x_{1}f\end{matrix}$
$\begin{matrix}\mathrm{Ind}_{\lambda_{2},\lambda_{0}}&\colon&\overbrace{\mathbf{Q}[x_{1},x_{2}]}^{=\mathrm{H}^{*}(V^{\lambda_{2}}/G^{\lambda_{2}})}&\rightarrow&\overbrace{\mathbf{Q}[x_{1},x_{2}]}^{=\mathrm{H}^{*}(V/G)}\\ &&f&\mapsto&x_{2}f\end{matrix}$
$\begin{matrix}\mathrm{Ind}_{\lambda_{3},\lambda_{0}}&\colon&\overbrace{\mathbf{Q}[x_{1},x_{2}]}^{=\mathrm{H}^{*}(V^{\lambda_{3}}/G^{\lambda_{3}})}&\rightarrow&\overbrace{\mathbf{Q}[x_{1},x_{2}]}^{\mathrm{H}^{*}(V/G)}\\ &&f&\mapsto&x_{1}x_{2}f\end{matrix}$
$\begin{matrix}\mathrm{Ind}_{\lambda_{3},\lambda_{1}}&\colon&\overbrace{\mathbf{Q}[x_{1},x_{2}]}^{=\mathrm{H}^{*}(V^{\lambda_{3}}/G^{\lambda_{3}})}&\rightarrow&\overbrace{\mathbf{Q}[x_{1},x_{2}]}^{=\mathrm{H}^{*}(V^{\lambda_{1}}/G^{\lambda_{1}})}\\ &&f&\mapsto&x_{2}f\end{matrix}$
$\begin{matrix}\mathrm{Ind}_{\lambda_{3},\lambda_{2}}&\colon&\overbrace{\mathbf{Q}[x_{1},x_{2}]}^{=\mathrm{H}^{*}(V^{\lambda_{3}}/G^{\lambda_{3}})}&\rightarrow&\overbrace{\mathbf{Q}[x_{1},x_{2}]}^{=\mathrm{H}^{*}(V^{\lambda_{2}}/G^{\lambda_{2}})}\\ &&f&\mapsto&x_{1}f\end{matrix}$
@@
$\mathcal{P}_{\lambda_{0}}=\mathbf{Q}\subseteq\mathbf{Q}[x_{1},x_{2}],\quad\mathcal{P}_{\lambda_{1}}=\mathbf{Q}\subseteq\mathbf{Q}[x_{1},x_{2}],\quad\mathcal{P}_{\lambda_{2}}=\mathbf{Q}\subseteq\mathbf{Q}[x_{1},x_{2}],\quad\mathcal{P}_{\lambda_{3}}=\mathbf{Q}\subseteq\mathbf{Q}[x_{1},x_{2}]\,.$
We have
$G_{\lambda_{0}}=1\subseteq(\mathbf{C}^{*})^{2},\quad G_{\lambda_{1}}=\mathbf{C}^{*}\times 1\subseteq(\mathbf{C}^{*})^{2},\quad G_{\lambda_{2}}=1\times\mathbf{C}^{*}\subseteq(\mathbf{C}^{*})^{2},\quad G_{\lambda_{3}}=(\mathbf{C}^{*})^{2}\,.$
The cohomological integrality isomorphism is
$\begin{matrix}\overbrace{\mathbf{Q}}^{=\mathcal{P}_{\lambda_{0}}}\oplus\overbrace{\mathbf{Q}[x_{1}]}^{=\mathcal{P}_{\lambda_{1}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda_{1}})}\oplus\overbrace{\mathbf{Q}[x_{2}]}^{=\mathcal{P}_{\lambda_{2}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda_{2}})}\oplus\overbrace{\mathbf{Q}[x_{1},x_{2}]}^{=\mathcal{P}_{\lambda_{3}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda_{3}})}&\rightarrow&\overbrace{\mathbf{Q}[x_{1},x_{2}]}^{=\mathrm{H}^{*}(V/G)}\\ (a_{0},f(x_{1}),g(x_{2}),h(x_{1},x_{2}))&\mapsto&a_{0}+x_{1}f(x_{1})+x_{2}g(x_{2})+x_{1}x_{2}h(x_{1},x_{2})\,.\end{matrix}$
@@ 6.2. $\mathrm{GL}_{2}(\mathbf{C})\curvearrowright\mathbf{C}^{2}\oplus(\mathbf{C}^{2})^{\vee}$
where the superscript $t$ indicates the transpose. We let $T\coloneqq(\mathbf{C}^{*})^{2}\subseteq\mathrm{GL}_{2}(\mathbf{C})$ be the standard torus. The Weyl group is $W=\mathfrak{S}_{2}$. We have $\mathscr{P}_{V}/W=\{\widetilde{\lambda_{0}},\widetilde{\lambda_{1}},\widetilde{\lambda_{2}},\widetilde{\lambda}_{3}\}$ where
@@
$\begin{matrix}\lambda_{1}&\colon&\mathbf{G}_{\mathrm{m}}&\rightarrow&(\mathbf{C}^{*})^{2}\\ &&v&\mapsto&(v^{-1},1)\end{matrix}$
$\begin{matrix}\lambda_{2}&\colon&\mathbf{G}_{\mathrm{m}}&\rightarrow&(\mathbf{C}^{*})^{2}\\ &&v&\mapsto&(v^{-1},v^{-2})\end{matrix}$
$\begin{matrix}\lambda_{3}&\colon&\mathbf{G}_{\mathrm{m}}&\rightarrow&(\mathbf{C}^{*})^{2}\\ &&v&\mapsto&(v^{-1},v^{-1})\end{matrix}$
Therefore, we have $V^{\lambda_{0}}=V$, $V^{\lambda_{1}}=(0\oplus\mathbf{C})\oplus(0\oplus\mathbf{C}^{\vee})$, $V^{\lambda_{2}}=\{0\}$, and $V^{\lambda_{3}}=\{0\}$. We have $G^{\lambda_{0}}=\mathrm{GL}_{2}(\mathbf{C}),G^{\lambda_{1}}=T$, $G^{\lambda_{2}}=T$ and $G^{\lambda_{3}}=G$. We can describe the induction maps as follows:
$\begin{matrix}\mathrm{Ind}_{\lambda_{1},\lambda_{0}}&\colon&\overbrace{\mathbf{Q}[x_{1},x_{2}]}^{=\mathrm{H}^{*}(V^{\lambda_{1}}/G^{\lambda_{1}})}&\rightarrow&\overbrace{\mathbf{Q}[x_{1},x_{2}]^{\mathfrak{S}_{2}}}^{=\mathrm{H}^{*}(V/G)}=\mathbf{Q}[x_{1}+x_{2},x_{1}x_{2}]\\ &&f&\mapsto&\frac{x_{1}}{x_{1}-x_{2}}f(x_{1},x_{2})+\frac{x_{2}}{x_{2}-x_{1}}f(x_{2},x_{1})\end{matrix}$
$\begin{matrix}\mathrm{Ind}_{\lambda_{2},\lambda_{0}}&\colon&\overbrace{\mathbf{Q}[x_{1},x_{2}]}^{=\mathrm{H}^{*}(V^{\lambda_{2}}/G^{\lambda_{2}})}&\rightarrow&\overbrace{\mathbf{Q}[x_{1},x_{2}]^{\mathfrak{S}_{2}}}^{=\mathrm{H}^{*}(V/G)}=\mathbf{Q}[x_{1}+x_{2},x_{1}x_{2}]\\ &&f&\mapsto&\frac{x_{1}x_{2}}{x_{1}-x_{2}}f(x_{1},x_{2})+\frac{x_{1}x_{2}}{x_{2}-x_{1}}f(x_{2},x_{1})\end{matrix}$
$\begin{matrix}\mathrm{Ind}_{\lambda_{2},\lambda_{1}}&\colon&\overbrace{\mathbf{Q}[x_{1},x_{2}]}^{=\mathrm{H}^{*}(V^{\lambda_{2}}/G^{\lambda_{2}})}&\rightarrow&\overbrace{\mathbf{Q}[x_{1},x_{2}]}^{\mathrm{H}^{*}(V^{\lambda_{1}}/G^{\lambda_{1}})}\\ &&f&\mapsto&x_{2}f\end{matrix}$
$\begin{matrix}\mathrm{Ind}_{\lambda_{2},\lambda_{3}}&\colon&\overbrace{\mathbf{Q}[x_{1},x_{2}]}^{=\mathrm{H}^{*}(V^{\lambda_{2}}/G^{\lambda_{2}})}&\rightarrow&\overbrace{\mathbf{Q}[x_{1}+x_{2},x_{1}x_{2}]}^{\mathrm{H}^{*}(V^{\lambda_{3}}/G^{\lambda_{3}})}\\ &&f&\mapsto&\frac{f(x_{1},x_{2})-f(x_{2},x_{1})}{x_{1}-x_{2}}\end{matrix}$
@@
The induction $\mathrm{Ind}_{\lambda_{2},\lambda_{3}}$ is surjective as any symmetric polynomial $f$ is the image by $\mathrm{Ind}_{\lambda_{2},\lambda_{3}}$ of $(x_{1}-x_{2})f$.
A direct sum complement of the image of $\mathrm{Ind}_{\lambda_{2},\lambda_{1}}$ is $\mathbf{Q}[x_{1}]=\mathcal{P}_{\lambda_{1}}\otimes\mathbf{Q}[x_{1}]$ by definition of $\mathcal{P}_{\lambda_{1}}$.
Therefore, we have $\mathcal{P}_{\lambda_{0}}=0$, $\mathcal{P}_{\lambda_{1}}=\mathbf{Q}$, $\mathcal{P}_{\lambda_{2}}=\mathbf{Q}$ and $\mathcal{P}_{\lambda_{3}}=0$. We have $G_{\lambda_{0}}=1,G_{\lambda_{1}}=\mathbf{C}^{*}\times 1$, $G_{\lambda_{2}}=T$ and $G_{\lambda_{3}}=G$. The cohomological integrality isomorphism reads
$\begin{matrix}\overbrace{\mathbf{Q}[x_{1}]}^{=\mathcal{P}_{\lambda_{1}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda_{1}})}\oplus\overbrace{\mathbf{Q}[x_{1},x_{2}]^{\mathrm{sgn}}}^{=(\mathcal{P}_{\lambda_{2}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda_{2}}))^{\varepsilon_{V,\lambda_{2}}}}&\rightarrow&\overbrace{\mathbf{Q}[x_{1},x_{2}]^{\mathfrak{S}_{2}}}^{=\mathrm{H}^{*}(V/G)}\\ (f(x_{1}),g(x_{1},x_{2}))&\mapsto&\frac{x_{1}f(x_{1})-x_{2}f(x_{2})}{x_{1}-x_{2}}+2x_{1}x_{2}\frac{g(x_{1},x_{2})}{x_{1}-x_{2}}\end{matrix}$
where $\mathbf{Q}[x_{1},x_{2}]^{\mathrm{sgn}}$ is the sign-isotypic component, since the kernel $\frac{x_{1}x_{2}}{x_{1}-x_{2}}$ changes sign when we exchange the variables $x_{1},x_{2}$, and so the character $\varepsilon_{V,\lambda_{2}}$ of $W_{\lambda_{2}}=\mathfrak{S}_{2}$ given by Proposition 4.6 is the sign.